OpenStudy (anonymous):

I need fast help! find y such that the distance between the points is 8. (5,1) (5,y)

jimthompson5910 (jim_thompson5910):

Use the formula \[\Large d = \sqrt{\left(x_{2}-x_{1}\right)^2+\left(y_{2}-y_{1}\right)^2}\]

OpenStudy (anonymous):

I did that and got 64=0 + (1-y)^2.idk what happened..

jimthompson5910 (jim_thompson5910):

In your case, d = 8, x1 = 5, x2 = 5, y1 = 1 and y2 = y So \[\Large d = \sqrt{\left(x_{2}-x_{1}\right)^2+\left(y_{2}-y_{1}\right)^2}\] becomes \[\Large 8 = \sqrt{\left(5-5\right)^2+\left(1-y\right)^2}\]

jimthompson5910 (jim_thompson5910):

Now that you know that 64=0 + (1-y)^2 You can get rid of that 0 to get 64 = (1-y)^2

jimthompson5910 (jim_thompson5910):

What's next?

OpenStudy (anonymous):

umm....

OpenStudy (anonymous):

sqrt?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

so 8=1-y? distance is 7?

jimthompson5910 (jim_thompson5910):

no, the given distance is 8

OpenStudy (anonymous):

what is the 7?

OpenStudy (anonymous):

-7

jimthompson5910 (jim_thompson5910):

8 = 1 - y 8 - 1 = -y 7 = -y -y = 7 y = -7

OpenStudy (anonymous):

So is that my answer

jimthompson5910 (jim_thompson5910):

So the distance between (5,1) (5,-7) is 8 units

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thanks so much

jimthompson5910 (jim_thompson5910):

sure thing